OpenStudy (anonymous):

V varies directly with the cube of r. V=36π and r =3 Please tell me how you do it

OpenStudy (akshay_budhkar):

what is that you want?

OpenStudy (akshay_budhkar):

the relation between V and r?

OpenStudy (anonymous):

k first I think

OpenStudy (anonymous):

So V = kr^3...

OpenStudy (anonymous):

yes but my friend told me that I have to put pi on 3 to the 3rd

OpenStudy (akshay_budhkar):

V = kr^3. you ok till here?

OpenStudy (anonymous):

Well u have V = 36pi and r = 3.

OpenStudy (anonymous):

yes

OpenStudy (akshay_budhkar):

goin ahead of estudier substitute to get 36pi=k(27) as 3cube is 27

OpenStudy (akshay_budhkar):

now divide by 27 on both the sides

OpenStudy (anonymous):

ohh so that's it.. I thought you're going to arrange it to subtract 36 to 27

OpenStudy (akshay_budhkar):

why are you subtracting? just divide to get 36pi/27=k...

OpenStudy (akshay_budhkar):

reduce it to 4/3(pi)

OpenStudy (anonymous):

yes I understood

OpenStudy (akshay_budhkar):

good.. :)

OpenStudy (anonymous):

Now I'm going to substitute everything thanks :D